OpenStudy (anonymous):

Use Laplace transforms to find a nontrivial solution to tx''+(3t−4)x'+3x=0, x(0)=0

OpenStudy (anonymous):

it also states to Require that x(1)=e^(−3) to find any arbitrary constant in your solution.

OpenStudy (anonymous):

@Tomas.A

OpenStudy (anonymous):

@Spacelimbus I think I almost got it

OpenStudy (anonymous):

as of now I got (s+3)X'(s)+6X(s)=0

OpenStudy (anonymous):

now we're supposed to X'(s)/X(s)=(s+3)/6

OpenStudy (anonymous):

no wait it's -3/(s+3)

OpenStudy (anonymous):

no i think -6/(s+3)

OpenStudy (anonymous):

never mind! i solved it! woooo